### v2

# LaMSUM: Creating Extractive Summaries of User Generated Content using LLMs

Abstract.
Large Language Models (LLMs) have demonstrated impressive performance across a wide range of NLP tasks, including summarization. LLMs inherently produce abstractive summaries by paraphrasing the original text, while the generation of extractive summaries – selecting specific subsets from the original text – remains largely unexplored. LLMs have a limited context window size, restricting the amount of data that can be processed at once. We tackle this challenge by introducing LaMSUM, a novel multi-level framework designed to generate extractive summaries from large collections of user-generated text using LLMs. LaMSUM integrates summarization with different voting methods to achieve robust summaries. Extensive evaluation using four popular LLMs (Llama 3, Mixtral, Gemini, GPT-4o) demonstrates that LaMSUM outperforms state-of-the-art extractive summarization methods. Overall, this work represents one of the first attempts to achieve extractive summarization by leveraging the power of LLMs, and is likely to spark further interest within the research community.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanyisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Brevity is the soul of wit.
— William Shakespeare, Hamlet, Act 2, Scene 2
Social media platforms like Facebook, X (formerly Twitter), and Reddit offer a medium for individuals to express their opinions and views on various subjects, leading to a diverse array of perspectives shared through social debates, critiques, and reviews  (Zhang et al., 2024b; Chhikara et al., 2023; Dash et al., 2019a; Bražinskas et al., 2020; Zhang et al., 2016). With a vast amount of data being generated by users online, there is a growing need for summarization algorithms providing a precise and concise summary, eliminating the need for users to sift through numerous posts or reviews. Based on the resulting summary, summarization algorithms can be categorized as ‘extractive’ and ‘abstractive’. In extractive summarization, the aim is to select a subset representative of the original text (Xu et al., 2020; Liu and Lapata, 2019; Zhong et al., 2020; Zhang et al., 2022, 2023a). In contrast, abstractive summarization aims to generate natural language summaries that capture the essence of the original text, often paraphrasing the content rather than directly extracting it (Lewis et al., 2020; Zhang et al., 2020).
Extractive summarization is a crucial task with a rich body of literature, widely applied in summarizing legal cases, news articles, lectures, clinical notes, and social media content (Bhattacharya et al., 2021; Kanwal and Rizzo, 2022; Mukherjee et al., 2020; Jia et al., 2020). For example, a Google search for a topic or hashtag displays a few tweets alongside the usual lists of websites and news articles. The selection of tweets displayed in search results is similar to an extractive summarization task, where only a subset of tweets is selected (Griffin, 2015). Additionally, the Library of Congress only stores a selection of tweets as part of its archive to optimize storage space (Congress, 2017). E-commerce platforms like Amazon display only a subset of reviews in the condensed view, such that the chosen reviews are representative of all the reviews submitted by the users. Such cases are instances of extractive summarization, where the target is to select the most relevant subset that can effectively summarize a topic or a product. When summarizing user generated data, it is crucial to quote the user’s exact words, making extractive summarization particularly valuable in such cases. In the process of summarization, each post or review is treated as a basic unit – often referred to as a textual unit.
In recent years, Large Language Models (LLMs) have exhibited high performance across various tasks, including summarization (Ouyang et al., 2022; Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024). Summaries generated by LLMs showcase high coherence and are overwhelmingly preferred by the human evaluators (Pu et al., 2023; Liu et al., 2024). However, using LLMs for extractive summarization has two significant limitations: (i) as generative models, LLMs naturally tend to perform abstractive summarization by paraphrasing rather than selecting the most relevant sentences; and (ii) due to the finite size of the context window, LLMs cannot handle long texts in a single input, underscoring the need for a method that allows for processing long text. Given the increasing reliance on LLMs for summarization, it is worth investigating the utility of LLMs for extractive summarization of large volumes of user-generated text.
To address these limitations, we present a novel framework LaMSUM (Large Language Model based Extractive SUMmarization) that integrates LLM-generated summaries with voting algorithms borrowed from Social Choice Theory (Brandt et al., 2016). Voting algorithms are traditionally employed in decision making processes to aggregate preferences and opinions of a large population; the prime example being the electoral system, where voters’ preferences for candidates are gathered to elect a winner that best represents the voters’ preferences. Voting systems can be categorized into single-winner and multi-winner types, depending on the number of winners. In a single-winner system, only one candidate is elected, whereas a multi-winner system allows for the selection of multiple candidates.
Extractive summarization can be viewed as a multi-winner election, where the input units act as candidates, and the units chosen for the summary are considered as the winners. A multi-winner voting system includes approval-based committee (ABC) voting, where voters approve a subset of candidates without ranking, and ranked choice voting, where voters rank candidates by preference. In LaMSUM, we utilize two approval based voting algorithms – Plurality Voting (Mudambi et al., 1996) and Proportional Voting (Lackner et al., 2023), and one ranked based voting algorithm – Borda Count (Emerson, 2013). Our judicial application of voting algorithms with a multi-level summarization framework ensures that LaMSUM outperforms the state-of-the-art fine-tuned summarization models. In summary, in this work, we make the following contributions:
  • We propose a novel framework LaMSUM which can effectively summarize large (having ¿30K tokens) collection of user generated content.
  • LaMSUM considers a multi-level summarization model that utilizes voting algorithms to combine outputs to generate robust summaries.
  • Analyse whether an ensemble model with multiple LLMs in LaMSUM can lead to improved outcomes across different voting methods.
To our knowledge, this is the first work to implement extractive summarization of large user-generated texts using LLMs by combining summarization with voting algorithms. We believe this work can spawn further research in this direction.
2. Background and Related Work
In this section, we review the relevant prior works that provide the foundation for our current research.
Figure 1. Current LLMs, by default, produce abstractive summaries. Llama3-70b-8192, despite specifically prompted for extractive summarization, generates abstractive summaries. This behavior underscores the need for a targeted approach to enable LLMs to effectively generate extractive summaries.
Text Summarization Algorithms
With the growing volume of online data, the demand for algorithms that automatically shorten and summarize texts is increasing. Automatic summarization can be approached in two ways: extractive and abstractive. In extractive summarization, a subset of input collection is selected for the summary based on their perceived quality and significance, aiming to represent a larger dataset with a concise sample. Abstractive summarization, on the other hand, generates natural language summaries that capture the most critical information from the original text. Both approaches seek to provide readers with a concise overview of the textual content. Over the years, many text summarization algorithms have been proposed in the literature; the reader can refer to (Gupta and Lehal, 2010; El-Kassas et al., 2021) for detailed surveys.
Extractive Summarization of User Generated Content
A large body of prior research has focused on summarizing long documents (Allahyari et al., 2017; Dong, 2018; Cajueiro et al., 2023). In fact, traditionally, summarization tasks have focused on summarizing a single document, such as a news article or a business report. In recent years, summarization has been increasingly applied on different types of user generated text (e.g., tweets, Facebook or Reddit posts) (Mackie et al., 2014; Inouye and Kalita, 2011; Jiang et al., 2021), where the task is to summarize short, independent posts written by many users. Several extractive summarization algorithms tailored specifically for user generated content have also been proposed (Sharifi et al., 2010; Xu et al., 2013; Dash et al., 2019b; Mukherjee et al., 2022; Rudra et al., 2018; Jones et al., 2023).
Large Language Models (LLMs) for Summarization
LLMs are now being extensively used for summarization (Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024). Multiple works have proposed few-shot learning frameworks for the abstractive summarization of news, documents, webpages, and generic texts (Zhang et al., 2023b; Tang et al., 2023a; Yang et al., 2023; Bražinskas et al., 2020; Laskar et al., 2023), but their primary focus remains on short documents that can fit in the LLM context window. Researchers have also observed that human evaluators are increasingly preferring LLM-generated summaries compared to other baselines (Zhang et al., 2024a; Wu et al., 2024; Goyal et al., 2023; Zhang et al., 2023c; Liu et al., 2024). Despite the advancements, recent studies have also uncovered factual inaccuracies and inconsistencies in LLM-generated summaries (Tang et al., 2024; Tam et al., 2023; Luo et al., 2023; Laban et al., 2023).
Figure 2. LaMSUM: Multi-level framework for extractive summarization of large user-generated text. Input set $\mathcal{T}$ (level 0) is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks each of size s. From each chunk a summary is produced of size $q$ (refer Figure 3), $q$ length summaries from $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks are merged to form the input for the next level i.e., level 1. Iteratively the same procedure is repeated till we obtain a summary of size $k$. We set $q=k$ to ensure our algorithm can effectively handle the worst-case scenario where all the textual units in the final summary may come from the same input chunk (Section 3.2).
Extractive Summarization through LLMs: The Current State By default, LLMs produce abstractive summaries, meaning that the summary text is distinct from the input text, even when it is instructed to do otherwise. To illustrate this, we present a small example in Figure 1. An LLM, when prompted, could clearly explain extractive summarization, yet, when we instructed it to perform extractive summarization on a set of 100 sentences, it fails to do so and instead generated an abstractive summary. Prior to our current work, only two studies attempted to perform similar tasks. Zhang et al. (2023b) attempted summarization of short news articles using GPT 3.5, while Chang et al. (2024) attempted abstractive summarization for book-length documents. However, both these approaches suffer from practical limitations such as lack of contextual dependencies in user generated text and the problem with positional bias in LLMs.
To the best of our knowledge, ours is the first attempt to perform extractive summarization on a large collection of user generated texts through LLMs, while tackling the challenge of positional bias. We describe our proposal in detail in the next section.
3. Generating Extractive Summaries through LLMs
In this section, we define the problem statement formally and introduce our novel summarization framework LaMSUM (Large Language Model based Extractive SUMmarization) that leverages LLMs to summarize large user-generated text.
3.1. Task Formulation
[…]
3.2. Multi-Level Summarization
LLMs have a limited context window, making it impossible to input large text collections all at once. While recent models like GPT-4 support context windows of up to 128k tokens, they still cannot accommodate book-length inputs within a single window. Consequently, the input must be divided into smaller chunks to perform the desired task (Chang et al., 2024). Thus, LaMSUM employs a multi-level framework for extractive summarization, enabling it to consider input data of any size (detailed in Figure 2)  [^1] Code is available at https://anonymous.4open.science/r/LaMSUM/ .
The set $\mathcal{T}$, which contains the original textual units, is provided as input at level 0 and is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ number of chunks of size $s$. From each chunk of size $s$, we generate a summary (discussed in Section 3.3) of size $q$ (where $q<s$), and repeat this process for all $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks.  [^2] Note that a chunk of size $s$ refers to a chunk containing $s$ textual units. Likewise, a summary of size $q$ indicates a summary of $q$ textual units. $|\mathcal{T}|$ denotes the number of textual units present in $\mathcal{T}$. We then merge all these $q$ length summaries obtained from level 0 to form an input for the next level i.e., level 1. We repeatedly perform this process until we obtain the final summary of length $k$. Note that the last chunk maybe less than $q$ in size, in such case we move all the textual units of the respective chunk to the next level (refer Algorithm 1).
An alternate strategy would be to divide the input $\mathcal{T}$ into $\frac{|\mathcal{T}|}{s}$ chunks each of size $s$ and from each chunk select $\frac{k\cdot s}{|\mathcal{T}|}$ sentences to be included in the summary. However, this approach assumes a uniform distribution of potential candidates across chunks that can be included in the final summary. In LaMSUM, we keep $q=k$ i.e., we extract $k$ textual units from each chunk eliminating the chance of missing any potential candidate (discussed in Section 4.4). In the worst-case scenario, all $k$ units in the final summary can come from a single chunk, and our algorithm can handle such cases effectively.
It is important to note that we are dealing with user-generated content, such as tweets, which lack contextual connections. Unlike book summarization, where chapters are interconnected and the context of previous chapters is crucial for summarizing the current one, tweets are generally standalone and contextually independent. Thus, our approach of independently deriving summaries from each chunk works well in our setup, as each textual unit operates independently of the others and there are no long-range dependencies.
Figure 3. Textual units (e.g., posts) in the input chunk are shuffled to account for the positional bias. $m$ different chunk variations are obtained through shuffling, which are subsequently summarized using LLMs. $m$ summaries are then aggregated by voting algorithms to get the final summary.
3.3. Summarizing a Chunk
Next, we discuss how LaMSUM summarizes a chunk (Algorithm 2) by tackling the positional bias in LLMs and leveraging voting algorithms drawn from Social Choice Theory (Brandt et al., 2016).
3.3.1. Tackling Positional Bias
Prior research  (Brown and Shokri, 2023; Zhang et al., 2023b; Jung et al., 2019; Wu et al., 2024) has highlighted that summarization using LLM is prone to positional bias, i.e., the sentences located in certain positions, such as the beginning of articles, are more likely to be considered in the summary. To address this issue and generate a robust summary, we create $m$ different variations by shuffling the textual units within the input chunk. This ensures that each unit has the opportunity to appear in different positions within the input text (refer Figure 3).
3.3.2. Zero-shot prompting
For each input chunk, we obtain $m$ different summaries (one for each variation) by prompting the LLM. We employ the following two prompts to obtain the summaries –
a
Select the most suitable units that summarize the input text.
Prompt: ”Input consists of ¡chunk_size¿ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that selects best ¡summary_length¿ sentences (subset) which summarizes the input. Think step by step and follow the instructions. ¡sentences¿”
b
Generate a ranked list in descending order of preference.
Prompt: ”Input consists of ¡chunk_size¿ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that outputs the sentences in the decreasing order of their relevance to be included in the summary. Think step by step and follow the instructions. ¡sentences¿ Remember that output should contain all the sentences in the decreasing order of their preference.”
Algorithm 1 Algorithm for multi-level summarization
Input: ${\mathcal{T},k,s,q,m}$
${}SS{\SS=\{\}}$triangleright $SS\SS$ stores the final summary
[…]
     $n_{chunks}=\lceil{\frac{|\mathcal{T}|}{s}}\rceil$triangleright number of chunks in set $\mathcal{T}$
     $L=\{\}$triangleright $L$ stores the results of a given level
     for $i\leftarrow 1$ to $n_{chunks}$ do
         $si=(i-1)*s$triangleright starting index of chunk
         $ei=i*s$triangleright ending index of the chunk
         if $i=n_{chunks}$ then $\triangleright$ if last chunk
              $ei=|\mathcal{T}|$triangleright ending index is equal to length of $\mathcal{T}$
         end if
         $width=ei-si$triangleright number of textual units in a chunk
         if $width<=q$ then $\triangleright$ if last chunk
              $L=L\cup t_{si}\cup t_{si+1}\cup...\cup t_{ei-1}$
$\triangleright$ add all textual units to the result
         else
              $L=L\cup\textsc{ChunkResult}(\mathcal{T},si,ei,q,m)$
$\triangleright$ add summary of each chunk to result $L$
         end if
     end for
     $\mathcal{T}=L$triangleright update the input $\mathcal{T}$ for the next level
[…]
end while
Output: $SS\SS$
Algorithm 2 Algorithm for summarization of a chunk
function ChunkResult($\mathcal{T},si,ei,q,m$)
     $X=\{\}$
     for $i\leftarrow 1$ to $m$ do $\triangleright$ for each variation of a chunk
         $V=\textsc{Shuffle}(\mathcal{T},si,ei,i)$triangleright shuffle with random state $i$
         $R=\textsc{LLM}(V,q)$triangleright obtain summary from the LLM
         $C=\textsc{Check}(R,\mathcal{T},si,ei)$triangleright output calibration
         $X.add(C)$
     end for
     return $\textsc{VOTING}(X,q)$triangleright voting to obtain the final summary
end function
 Original Post LLM Modified Output
  My allergies be acting up &amp; Claritin be letting them &gt;,&lt; My allergies be acting up & Claritin be letting them.
I spent the extra money to have fast relief, but instead I am SUFFERING #gimmemymoneyback I spent the extra money to have fast relief, but instead I am SUFFER
Table 1. Examples illustrating that LLMs when selecting textual units for summarization, often demonstrate a propensity to alter certain words or introduce new ones.
3.3.3. Output Calibration
LLMs may alter certain words from the input text while generating extractive summaries, as shown in Table 1. Thus, we perform additional checks to ensure that the textual units selected in the summary are indeed a subset of $\mathcal{T}$. If the post selected by the LLM (say $x$) is not present in the original text $\mathcal{T}$, we identify the post with the closest resemblance to $x$ by computing the edit distance (Ristad and Yianilos, 1998). LLMs may also hallucinate, generating new sentences rather than selecting units from the input. In such instances, the edit distance between the generated unit $x$ and all the original textual units would be high. We discard these sentences from the output.
3.4. Reimagining Summarization as an Election
As mentioned earlier, for a given chunk, we obtain $m$ summaries – one for each variation. We imagine the process of creating the final summary from these $m$ summaries to be a multi-winner election, where the textual units in $m$ summaries correspond to ballots (candidates) and the role of the voting algorithm is to pick $q$ winners. We employ three different voting methods, namely Plurality Voting (Mudambi et al., 1996), Proportional Approval Voting (PAV) (Lackner et al., 2023) and Ranked Choice Voting (Emerson, 2013) to determine the final summary. Due to the varying input requirements of different voting methods, changing both the prompting approach and the output generated by the LLM becomes imperative.
Plurality voting and proportional voting are approval-based voting methods where voters can select multiple candidates they approve of without indicating a specific preference order. In multi-winner plurality voting (also known as block voting), each voter casts multiple votes and the candidates are selected based on the number of votes polled. In the context of summarization, a textual unit is treated as a candidate, and the LLM acts as the voter. We select the textual units in the decreasing order of the votes polled, till we obtain a summary of size $q$. PAV evaluates the satisfaction of each voter in the election outcome. A voter’s satisfaction is measured based on – amongst the number of candidates they voted for, how many are selected in the election. In the realm of summarization, PAV selects the textual units based on the amount of support each unit receives in $m$ summaries. Since both plurality and proportional are approval-based voting algorithms, the units are either approved or disapproved by the underlying LLM, with no explicit ranking or preference order. In this case, we prompt the LLM to select the best ¡$q$¿ sentences that summarize the input text as shown in Section 3.3.2 a.
On the other hand, ranked choice voting entails assigning a score to each textual unit and subsequently selecting the highest-scoring units for inclusion in the summary. For ranked voting, we use the Borda count, a positional voting algorithm (Emerson, 2013). In the Borda method, each candidate is assigned points corresponding to the number of candidates ranked below them: the lowest-ranked candidate receives 0 points, the next lowest gets 1 point, and so forth. The candidates with the highest aggregate points are declared as the winners. In ranked voting, we prompt the LLM to output sentences in descending order of their suitability for the summary as discussed in Section 3.3.2 b.
It is important to note that the prompting technique and the output generated by LLM vary for different voting methods. In approval voting the output from LLM is a list of $q$ textual units that LLM finds best suited to be included in the summary. Whereas in ranked choice voting, the output from LLM is a list of the same length as input i.e. $s$ with all the units sorted in decreasing order of their preference towards the summary, and Borda Count (Emerson, 2013) is used to identify the top $q$ textual units. In the next section, we highlight how the voting-based summarization schemes outperform the Vanilla setup, which does not use voting.
4. Experimental Setup
      Parameters Claritin US-Election MeToo
      #TextualUnits ($|\mathcal{T}|$) 3998 2107 483
#InputWords 53609 35522 16737
AV #CTU ($s$) 200 150 75
#CSTU ($q$) 100 100 50
RV #CTU ($s$) 40 40 40
#CSTU ($q$) 20 20 20
#SummaryTU ($k$) 100 100 50
Table 2. Input parameters used for the proposed framework LaMSUM. #TextualUnits is the number of textual units in the input set i.e. $|\mathcal{T}|$. #InputWords represents the number of words present in the input set $\mathcal{T}$. #CTU is the number of textual units in a chunk i.e. chunk size represented as $s$. #CSTU is the number of textual units in the chunk summary represented as $q$. #SummaryTU is the number of textual units present in the final summary i.e. $k$. AV and RV represents Approval Voting and Ranked Voting respectively.
4.1. Dataset
Our experiments are conducted on three publicly available datasets, consisting of crowd-sourced data from X, listed in Table 2 (Dash et al., 2019a). The Claritin dataset contains 3,998 tweets about the benefits and the side-effects of the anti-allergic drug Claritin. US-Election dataset contains 2,107 tweets from 2016 US Presidential Election where people support and attack different political parties. Me-Too dataset includes 483 tweets from the October 2018 MeToo movement, where individuals recount the harassment cases they experienced. [^3] Dataset: https://github.com/ad93/FairSumm/tree/master/Dataset During pre-processing, we remove all web links and duplicate entries from the datasets. The datasets also include gold-standard summaries (reference summaries) which are human-generated i.e., the textual units that are strong candidates for inclusion in the summary are selected by the humans. Claritin and US-Election dataset, each have three gold standard summaries, comprising 100 textual units each. Me-Too dataset has two gold-standard summaries with 50 textual units each.
4.2. Large Language Models (LLMs)
LLMs are characterized by their extensive parameter sizes and remarkable learning abilities (Zhao et al., 2023; Chang et al., 2023). In our work, we utilize three open-source LLMs and one proprietary LLM to conduct experiments: llama3-70b-8192 from Meta (Touvron et al., 2023), mixtral-8x7b-32768 from Mistral AI (Jiang et al., 2024), gemini-1.0-pro from Google (Gemini Team, 2023) and
gpt-4o-mini-2024-07-18 from OpenAI (OpenAI, 2024). Across all experiments, we keep temperature, top probability and output tokens as 0, 0.8 and 8192 respectively.
4.3. Evaluation Metric
For evaluating the quality of summaries generated by LaMSUM, we report ROUGE-1, ROUGE-2, and ROUGE-Lsum scores (Lin, 2004). ROUGE-1, ROUGE-2 and ROUGE-L respectively evaluate the overlap of unigrams, bigrams and longest common subsequence between the generated summary and the reference summary. ROUGE-Lsum is more suitable for extractive summarization, as it applies ROUGE-L at sentence level and then aggregates all the results to obtain the final score.
4.4. LaMSUM Input Parameters
Input parameters for LaMSUM (Algorithm 1), such as $|\mathcal{T}|$ (total number of textual units in the set), $s$ (chunk size) and $k$ (length of summary) are listed in Table 2 for different voting algorithms and datasets. The value of $m$ (number of shuffling) for all the datasets was set to 5.
If $q\in[k,s)$, our proposed method can handle worst case scenario where all the textual units in the final summary may originate from a single chunk of level 0. As $q$ approaches $s$, more levels are required to converge to the final summary. The optimal value of $q$ that can handle worst case and also reduce the number of levels in multi-level summarization is $k$, thus we keep $q=k$ for experiments with approval voting. For instance, if $s$ is 200 and $q$ is 100, this indicates that only 50% of the units from each chunk advance to the next level.
In ranked voting algorithm, we maintain smaller value for chunk size ($s$) to ensure that the LLMs output, which is of size $s$, fits within the context window. Additionally, as chunk size increases, LLM often does not output all the sentences, instead produce generalized statements like “similarly for other sentences we find the rank”. Therefore, it is essential to keep the chunk size smaller. For ranked voting, we set $s$ and $q$ to 40 and 20 respectively, upholding the selection ratio of 50% at each level.
5. Experimental Evaluation
In this section, we present the empirical comparison of LaMSUM with competent baseline models and voting algorithms across datasets (refer Table 3).
                      Models Claritin US-Election         MeToo
R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
                     ClusterRank 50.08 14.17 48.64 55.18 13.09 53.51 55.81 55.81 52.48
DSDR 28.01 8.40 27.45 48.04 9.53 46.44 55.03 17.30 52.04
LexRank 45.04 19.71 44.74 42.63 10.78 41.64 42.70 11.32 40.91
SummBasic 58.25 19.29 56.76 55.36 12.43 53.94 58.23 21.53 56.07
GPT2 61.61 23.58 60.74 55.86 15.07 54.81 40.63 11.24 38.86
BERT 57.30 22.37 56.21 55.89 15.44 55.00 45.72 10.76 43.50
XLNET 55.52 21.37 54.75 56.48 15.72 55.41 36.58 08.50 34.48
BERTSUM (Liu and Lapata, 2019) 57.87 22.75 55.96 59.00 17.51 57.41 57.11 23.08 54.84
 Vanilla Llama3 58.07 21.86 56.73 56.12 12.93 54.71 51.84 17.84 49.47
Mixtral 57.80 20.14 55.81 57.11 14.20 55.79 49.63 11.29 46.16
Gemini 49.70 18.33 48.78 51.00 12.36 49.19 45.51 8.95 42.51
GPT-4o mini 62.93 24.82 61.08 55.06 15.01 53.88 36.30 7.05 34.54
  Plurality Voting Llama3 61.28 23.79 59.54 60.11 18.26 58.99 55.51 18.79 52.85
Mixtral 59.13 22.90 57.40 59.55 16.01 58.12 55.41 17.62 52.63
Gemini 55.43 19.32 53.35 58.83 16.25 57.54 57.95 21.47 54.87
GPT-4o mini 64.20 26.71 62.66 58.15 15.91 56.78 54.55 19.66 51.9
Proportional Voting Llama3 61.31 23.36 59.73 58.90 15.99 57.55 58.14 14.99 55.46
Mixtral 60.30 24.03 58.58 58.46 15.74 57.28 57.36 22.97 54.82
Gemini 61.77 19.83 59.81 57.92 15.85 56.49 54.02 13.95 50.90
GPT-4o mini 64.13 26.75 62.30 58.68 19.08 57.62 48.55 13.23 45.35
 Borda Count Llama3 51.87 15.93 49.61 51.95 13.49 50.60 53.35 16.99 50.38
Mixtral 56.07 22.28 54.65 48.71 13.37 47.48 53.24 18.80 50.96
Gemini 51.45 16.92 49.39 54.01 13.38 52.56 50.13 21.45 47.96
GPT-4o mini 58.01 18.08 56.05 53.10 15.94 51.63 50.10 17.82 47.75
Table 3. Table showing metric scores from different models for various datasets. Here, R1 = ROUGE-1 Score, R2 = ROUGE-2 Score, RLSum = ROUGE-LSum Score. The best value per dataset is shown in bold and clearly Approval voting outperforms all the other methods across all the evaluation measures.
5.1. Baseline Comparison
We compare LaMSUM with the pre-neural models (ClusterRank (Garg et al., 2009), DSDR (He et al., 2012), LexRank (Erkan and Radev, 2004), SummBasic (Nenkova and Vanderwende, 2005)), transformer based models (GPT2 (Radford et al., 2019), BERT (Miller, 2019), XLNET (Yang et al., 2019)) and with fine-tuned BERTSUM (Liu and Lapata, 2019) model. As presented in Table 3 it is observed that LaMSUM outperform the state-of-the-art summarization models. Earlier work (Zhang et al., 2023b) reported that the ChatGPT model achieves lower ROUGE scores on CNN/DM and XSum dataset. But our results demonstrate that our proposed framework LaMSUM performs significantly better than other fine-tuned models for large user-generated text.
Results from the LLM exhibit variability when executed multiple times over the same input. We conducted experiments with LLMs for five iterations for the same input. Table 3 contains the maximum ROUGE score obtained from these five iterations and Table 6 (in supplementary material) displays the variance in the ROUGE scores when executed for five times.
Algorithm 3 Algorithm for summarization of a chunk in Vanilla LLM
function ChunkResult($\mathcal{T},si,ei,q,m$)
     $R=\textsc{LLM}(\mathcal{T},si,ei,q)$triangleright $q$ textual units from $\mathcal{T}\in[si,ei]$
     $C=\textsc{Check}(R,\mathcal{T},si,ei)$triangleright output calibration
     return $C$
end function
5.2. Vanilla LLM vs. LaMSUM
Our proposed framework, LaMSUM, ensures robust summary generation by shuffling and employing a voting algorithm to select the best textual units for the summary. It is crucial to compare LaMSUM with a multi-level LLM that does not use shuffling and voting, which we call Vanilla LLM. Algorithm 3 outlines the steps used by vanilla LLM to find the chunk summary. Table 3 demonstrates that the vanilla multi-level LLM has lower ROUGE scores for each LLM compared to the proposed framework LaMSUM, indicating that shuffling and voting enhances the performance.
5.3. Which Voting Algorithm Holds the Lead?
We experimented with three voting algorithms, two approval-based and one ranked-based. Experimental results indicate that LLMs with approval voting perform the best compared to the ranked voting algorithm. We hypothesized that rank-based voting would yield better results, as it makes more informed decisions about the potential sentences to be included in the summary. Contrary to our expectations, rank-based algorithms performed even worse than neural and transformer-based models. This can be attributed to multiple factors: (i) LLMs hallucinate and output sentences in the same or in the reverse order as they were in the input. (ii) Occasionally, LLMs do not output all the sentences from the input, resulting in the padding of left-out sentences towards the end of the list, which disturbs the ranking and potentially affects the result. To overcome these problems, we kept the chunk size low as discussed in Section 4.4, but the results still did not surpass those of the approval-based voting algorithm.
Takeaway: LLMs, when prompted to select sentences that can summarize the input, perform better than when tasked to rank the sentences in the order of their preference towards the summary.
5.4. What Fails to Deliver Results?
To ensure extractive summarization, we tested an additional approach – each sentence is tagged with a sentence number, LLM is prompted to select the best $q$ sentences and output only the sentence numbers of the best $q$ sentences. Thereafter, the sentences corresponding to the sentence numbers can be retrieved. For instance, if $s$ is 200 and $q$ is 100, the task is to output the sentence numbers of the best 100 sentences from a pool of 200 sentences. In such cases, LLMs hallucinate and provide an output consisting of either all the odd number sentences or all the even number sentences.
Takeaway: For extractive summarization, relying solely on indexes may result in hallucination, underscoring the importance of emitting the input content and not the numbers.
6. Ensembling of LLMs in LaMSUM
Figure 4. Ensembling applied at chunk using all the four models. $q$ length summaries are obtained from each chunk through different LLMs using the shuffling method discussed in Section 3.3. All the summaries are ensembled through voting algorithms to achieve the final summary of length $q$. This final $q$ length summary is thereafter forwarded to the next level.
 Ensemble Voting Method R1 R2 RLSum
 All 4 models Plurality 54.16 20.79 53.00
Proportional 54.70 19.95 53.14
Ranked 51.78 16.97 50.24
Best 3 models Plurality 58.88 21.74 57.60
Proportional 58.04 22.78 56.71
Ranked 57.22 19.51 55.33
Weak 3 models Plurality 50.64 18.82 49.31
Proportional 52.92 19.23 51.07
Ranked 47.50 16.74 46.37
Table 4. Results obtained through ensembling of LLMs in LaMSUM for Claritin dataset. We compare 3 different cases: i) ensemble of all 4 models – Llama, Mixtral, Gemini and GPT-4o mini ii). ensemble of best 3 models – Llama, Mixtral and GPT-4o mini iii). ensemble of weak 3 models – Llama, Mixtral and Gemini.
Different LLMs can be considered as distinct experts, each offering their vote or opinion about the textual unit to be included in the summary. In LaMSUM, each expert (LLM) operated independently, determining which sentences were worthy of inclusion. We hypothesize that using an ensemble of LLMs in LaMSUM, where these experts collaborate and reach a consensus, has potential to improve performance of the overall system (Verga et al., 2024; Schoenegger et al., 2024).
We apply ensembling at each chunk as shown in Figure 4. For a given chunk $i$ containing $s$ textual units, the goal is to produce a final summary of size $q$ by gathering opinions of all the experts. Ensembling takes place in two stages. Stage 1 - For each chunk we obtain $q$ length summary through different LLMs by using shuffling and voting method as demonstrated in Section 3.3. Stage 2 - To ensemble the $q$ length summaries from different LLMs, we again leverage the voting methods. In stage 2 we make use of the same voting method that was used in stage 1 to maintain consistency. The summary generated from the ensembling process is treated as the final summary for chunk $i$ and is carried forward to the next level.
As shown in Table 3, for Claritin dataset, GPT-4o mini outperforms the other models, while Gemini delivers weaker results. We evaluate three different scenarios based on the number of LLMs used in the ensembling process: i) all four models – Llama, Mixtral, Gemini and GPT-4o mini, ii) the top three models – Llama, Mixtral and GPT-4o mini, iii) the three weaker models – Llama, Mixtral and Gemini. Table 4 shows the results of ensembling in LaMSUM for different scenarios and voting methods.
In the approval-based voting algorithm, using all four models or just the top three models for ensembling doesn’t significantly improve performance; the resulting performance after ensembling remains close to that of the weakest model in each case. However, ensembling with the ranked-based voting algorithm yielded interesting results. When ranked voting is applied to individual LLMs, it delivers the worst performance compared to other voting algorithms, as shown in Table 3. When using ranked voting with all four models in the ensembled LaMSUM, the performance is close to that of the weakest model – Gemini (49.39) ¡ Llama (49.61) ¡ Ensemble (50.24) ¡ Mixtral (54.65) ¡ GPT-4o mini (56.05). However, when considering the ensembled LaMSUM with the top three models – Gemini (49.39) ¡ Llama (49.61) ¡ Mixtral (54.65) ¡ Ensemble (55.33) ¡ GPT-4o mini(56.05), ensemble outperforms Mixtral, and the ROUGE score increases from 50.24 to 55.33. While the ensembled LaMSUM with the top three models does not surpass GPT-4o mini, it does enhance the performance of the overall setup.
Takeaway: Ensemble of LLMs in LaMSUM yield good results with rank-based voting method, and can help improve the performance of the overall setup if we select models for ensembling that perform well individually in LaMSUM. Including all models in the ensemble could result in the weakest model becoming the bottleneck.
7. Does Fine-Tuning Help?
The performance of LLMs on specific downstream tasks, such as extractive summarization, can be suboptimal when not explicitly fine-tuned. LLMs tend to generate erroneous information while sounding persuasive and assured (Bang et al., 2023; Xu et al., 2024). In Section 2, we highlighted instances where LLMs exhibit nonplussed responses by generating abstractive summaries despite being prompted for extractive summarization. They also alter some words, so we perform output calibration in LaMSUM. To address these challenges, we explore the efficacy of instruction fine-tuning through explicit instructions (Wei et al., 2022) and compare it with LaMSUM.
We utilize the Claritin dataset for instruction fine-tuning as it has the highest number of tweets among all the three datasets. It is divided into training (60%), validation (10%), and test (30%) set. The US-Election and MeToo datasets are retained as additional test sets for evaluation. Instruction fine-tuning involves three key components: i) Instruction: The prompts used for zero-shot prompting, as detailed in Section 3.3.2. ii) Input: Set of input text $\mathcal{T}$ to be summarized. iii) Output: The reference summary comprises sentences extracted from $\mathcal{T}$.
We fine-tune an open-source LLM, llama3.1-8b from Meta (Dubey and et al., 2024), using a 15 GB Tesla T4 GPU. During the fine-tuning phase, we set the learning rate, warmup ratio, and batch size to 2e-4, 0.0, and 2, respectively. For this process, we employ PEFT (Parameter Efficient Fine Tuning), which involves freezing the layers of the pre-trained model and only fine-tuning the last few layers specific to the downstream task. We follow the QLoRA (Dettmers et al., 2023) method in PEFT training. QLoRA is an optimized variant of LoRA (Hu et al., 2021) that reduces the precision of weight parameters to 4-bit. This reduction in precision decreases the model size, which is advantageous in scenarios with limited memory available for fine-tuning.
Table 5 shows the results of the fine-tuned llama3.1-8b model. We observed that the outputs are robust and adhere closely to the instructions provided. The generated summaries contain the exact sentences extracted from the input text, but the performance of the fine-tuned model is inferior compared to LaMSUM.
 Dataset R1 R2 RLSum
 Claritin 18.60 3.11 17.14
US-Election 26.55 14.81 25.05
MeToo 28.94 10.81 24.80
Table 5. Metric scores from fine-tuned llama3.1-8b for various datasets.
Takeaway: Fine-tuning an LLM with the appropriate instructions, input, and output can facilitate extractive summarization. But fine-tuning requires a substantial amount of data with gold-standard summaries. With less data, the model may overfit, resulting in poor performance, as shown in the Table 5. Additionally, a model fine-tuned on one dataset may not perform well on another dataset. Our proposed LaMSUM addresses these issues by being effective when the input data is large enough to exceed a single context window but small enough to be used for fine-tuning.
8. Concluding Discussion
This work marks an early attempt to achieve extractive summarization of large user-generated text that exceeds a single context window using zero-shot learning. The proposed multi-level framework LaMSUM leverages approval based and ranked based voting algorithms to generate robust summaries. Experiments conducted on three distinct crowd-sourced datasets demonstrated the efficacy of LaMSUM, as it outperformed the results achieved by state-of-the-art fine-tuned models. Ensembling of LLMs in our proposed framework can enhance the performance of weaker models.
Note that there can be a concern regarding the potential data leakage, as the experiments involve newer LLMs that may have been exposed to the experimented datasets during their pre-training phase. In Section 5.2 we showcased that the vanilla LLM, which also includes LLM underperformed, whereas our proposed framework which generates robust summaries yielded good results. This highlights the efficacy of our model, even when it is exposed to data leakage.
Limitation
Our proposed framework, LaMSUM, very well handles text of any length, conditioned on the fact that the final summary fits within a single context window. Some modifications to LaMSUM may be necessary when the output summary exceeds the size of a single context window. We focused only on the user-generated text where each textual unit is independent of the others; future research could extend the framework to summarization tasks involving contextual dependency such as book summarization.
Ethical Considerations
Our research focuses on using LLMs to produce extractive summaries for user-generated text. Given the recent rise of LLMs and the growing interest in applying them across different research fields, we believe this research direction can help unveil the potential of LLMs. Below, we outline concerns that need to be considered and addressed in this research area:
  • Bias and discrimination: LLMs are biased towards their training data, which can lead them to favor certain textual units when creating summaries.
  • Transparency and accountability: LLMs are black boxes, with an opaque decision-making process, making it difficult to discern how or why specific textual units are chosen for summarization. This lack of transparency can create ethical issues, especially in scenarios where clear explanations for decisions are needed.
  • Environmental impact: The training and operation of LLMs require substantial computational resources, which contribute to increased energy consumption and may have a negative environmental impact.
While LLMs can produce high-quality summaries, their use must be approached with careful consideration of potential ethical implications.